It has a wire that is 3mm in diameter and 150m in length. Its resistance is 3Ω at 20 ºC. The resistivity of the wire.
Answer #1

given
d = 3 mm = 0.003 m
L = 150 m
R = 3 ohms

rho = ?

radius of the wire, r = d/2 = 0.003/2 = 0.0015 m
cross sectional area of the wire, A = pi*r^2

= pi*0.0015^2

= 7.07*10^-6 m^2

we know, R = rho*L/A

rho = R*A/L

= 3*7.07*10^-6/150

= 1.414*10^-7 ohm.m <<<<<<<----------------Answer